OpenStudy (anonymous):

5. LeAnne leaves town traveling at an average speed of 54 mph. After 2 hours, Bart leaves town traveling in the same direction at an average speed of 74 mph. Which of the following equations could be used to represent the distance between LeAnne and Bart after x hours? (Let x represent the time in hours that Bart has been traveling and y represent the distance between LeAnne and Bart.) y = 108 - 20x y = 108 - 74x y = 108 + 20x y = 20x

OpenStudy (anonymous):

at x = 2 we should have 108 since after 2 hours Bart leaves and LeAnne already traveled 54*2 miles now every hour the distance between then is decreasing by amount of 74-54 miles since its decreasing we have y=108-20x

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

yw:)

OpenStudy (anonymous):

8. Reid's Hardware discounts all riding lawnmowers 8% to customers paying in cash. If Trey paid $1,175.21 in cash for a riding lawnmower, which of the following equations can be used to determine the original price of the lawnmower? (Let x represent the original price of the lawnmower and y represent the discounted price.) y = 1.08x y = 0.92x y = x - 8x y = 1.8x

OpenStudy (anonymous):

@Coolsector ?

OpenStudy (anonymous):

if x is the original price then 0.08x is 8% of x so y = x - 0.08x y = 0.92x